OpenStudy (anonymous):

can you show me step by step how to put these two two equations into y=mx+b form? : 3x + y = 9 6x + 2y = 18

OpenStudy (anonymous):

3x + y = 9 0 + y = -3x + 9.... .anytime you want to get rid of a term bring it to the other side as a neg.

OpenStudy (anonymous):

6x + 2y = 18 0 + 2y = -6x + 18 now divide both sides by 2 y = -3x + 9

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

just use the formula of it. transfer the y to the left and transfer the 3x next to it but change the sign because your transposing it and then transfer last the 9 but do not change the positive sign of it because it's still positive. that's all i know, sorry for the long reply.